# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ 3 Kernel splitting I: helicity streams, color sum
Physics process $g\!\rightarrow\!t\bar{t} gg$$\rightarrow\!t\bar{t}g gg$$\rightarrow\!t\bar{t}gg gg$$\rightarrow\!t\bar{t}ggg gg$$\rightarrow\!t\bar{t}gggg gg$$g\!\rightarrow\!t\bar{t}ggggg$
Feynman diagrams 3 16 123 1240 15495 231280
SDE channels 3 15 105 945 N/A N/A
Leading colors 2 6 24 120 720 5040
Color matrix 2 x 2 6 x 6 24 x 24 120 x 120 720 x 720 5040 x 5040
Helicities 16 32 64 128 256 512
Wave functions 5 12 26 121 750 5834
Code generation codebase ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p1
Code generation mode mad mad mad mad sa sa
Code generation d.p.g. any any any any $\leq$100 $\leq$1000
Code generation d.p.f. any any any any $\geq$100 $\geq$1000
Code generation time 2s 2s 4s 32s 13m 30h
Feynman diagram groups 1 1 1 1 155 232
Feynman diagram files 1 1 1 1 155 232
LOC (color_sum.cc) 438 442 460 556 1156 5.4k
LOC (diagramsXYZ.cc) 135 307 1986 27k 652k 18.7M
Size (color_sum.cc) 24 kB 24 kB 28 kB 92 kB 2.6 MB 137 MB
Size (diagramsXYZ.cc) 8 kB 16 kB 84 kB 1.0 MB 28.2 MB 721 MB
Build time (five C++ modes) 6s 6s 6s 23s 3m 55m
Build time (CUDA DCDIAG=0) 11s 11s 14s 6m 9m failed
Build time (CUDA DCDIAG=1) 13s 13s 15s 3m 6m failed
@@
Table 2: Comparison of the computational complexity of the six physics processes considered in the tests described in this paper. The rows represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. of the diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for these six processes, this coincides with the number of good helicities with non-zero contributions to MEs); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the codebase used for code generation; (9) the code generation mode, “mad” for full generation including Fortran madevent and “sa” for standalone only; (9-10) the maximum number of diagrams per group (dpg) and the minimum number of diagrams per file (dpf) during code generation; (12) the code generation time on a typical CPU; (13-14) the resulting numbers of Feynman diagram groups and source code files; (15-17) the total number of lines of code and total size in the CUDAPP generated color_sum.cc and the sum of all diagramsXYZ.cc files; (19-21) the time for a parallel (-j) build of the software with a cold cache (CCACHE_RECACHE=1), either for all five C++ modes on an Intel Xeon Gold 6326, or for CUDA in each of two DCDIAG modes on an AMD EPYC 7313. All of these numbers are needed to describe the complexity of the ME calculation in a single standalone (or madevent) application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used, as of the ihel1 software. The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of ihel2, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel as of ihel2, and optionally via BLAS as of ihel3). The number of wavefunctions is particularly relevant to the ihel4-ihel6 software versions, which handle their storage and retrieval in GPU registers and/or GPU global memory in subtly different ways.
@@
As I briefly hinted above, one of the main aims of the new developments presented in this section is the fact that the ME calculation throughput (in MEs/s) on a GPU is generally quite low for small grids i.e. small values of nevt, and only reaches a peak plateau for relatively large grids, i.e. large values of nevt. This is an issue that exists in CUDACPP since the very beginning of our developments: in Fig. 5 of our vCHEP2021 proceedings bib:vchep2021 ), for instance, we had shown that the ME throughput for the $gg\!\rightarrow\!t\bar{t}gg$ process only ramps up significantly with at least 16k events in the grid, and reaches the peak plateau even later, with 128k events in the grid. In the same paper, we had also already suggested that higher throughputs, possibly because of lower “register pressure”, might be achieved by splitting monolithic sigmaKin kernel into smaller kernels: two specific ideas that we had mentioned, in particular, were the possible use of different GPU threads to process different helicities in the same event, or the possible use of CUDA graphs to orchestrate a much larger number of smaller GPU kernels. I will come back to both of these ideas in the following.
@@
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this paper will mainly consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers prepared for the HEP-SCORE benchmarking project bib:bmk ; it would be useful to repeat similar tests in the future, when updated HEP-SCORE containers are prepared using more recent versions of CUDACPP, possibly including those presented in this paper.
@@
In the following subsections, I will describe the first batch of three sets of kernel splitting changes that I developed, and which have already been integrated in a new production release of CUDACPP: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading. As shown in Fig. 1, these developments are sequential, i.e. ihel3 includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine for these three scenarios is represented in Fig. 3, where it is also compared to the situation before any kernel splitting, ihel0. Note that I will not provide new pseudo-code listings for these kernel splitting developments, but in some cases I will refer to the pseudo-code for version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, as well as for those in my second batch of changes described in the next section, are given in Fig. 4 for an NVidia V100 GPU at CERN, in Fig. 5 for an AMD MI200 GPU at LUMI, and in Fig. 6 for an Intel Xeon Silver 4216 CPU, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.
@@
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) cuBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, cuBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, cuBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
@@
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU (at LUMI). Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) hipBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, hipBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, hipBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
@@
Figure 6: Throughputs (ME/s) as a function of CUDACPP SIMD build mode for an Intel Xeon Silver 4216 CPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel3p1) release v1.01.01; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, using code generated with a single diagram group.
@@ 3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
Figure 7: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 version of the software. Lower is better. The 3 plots correspond to 3 floating point precisions. Measurements performed on a node equipped with an NVidia V100 GPU and an Intel Xeon Silver 4216 CPU (which supports AVX512 but gives better results for CUDACPP in the “512y” mode than in the “512z” mode because it has only one FMA unit). The results shown for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
@@
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. The achieved performance is represented by the “ihel3” solid-red curve in Fig. 4. As mentioned above, these results are indistinguishable from those of ihel2, which I therefore decided not to show this explicitly to simplify the plot; the comparison, however, is available in my previous preprint of this article bib:preprintV1 .
@@
    i.e. as the sum of two quadratic forms $V^{t}(C)V$, where the real vector $V$ may represent either the real part $A$ or the imaginary part $B$ of the complex dual amplitude vector $J$. Each calculation involves two steps. In mixed or float precision, for instance, the vector $(C)V$ is computed using cublasSgemm, while its dot product with $V^{t}$ is then computed using cublasSgemmStridedBatched. These two functions, as well as their double-precision and their HIP counterparts, are called through their abstractions via #define directives, in the header-only approach described in Ref. bib:chep2024 . The intermediate results of the first calculation, $(C)V$, are stored in GPU global memory using an additional buffer allocated in the MEK component (the allocation is done at runtime after determining the number of “good” helicities in each physics process, since the amount of memory allocated is proportional to the number of good helicities). As shown in the bottom-right diagram in Fig. 3, the BLAS calculations for different helicities are performed in separate GPU streams: technically, many BLAS handles are used, each associated to a different stream. Since the performance of the BLAS implementation of the color sum is generally worse than that using kernels, as discussed below, this is only available as an option, which must be explicitly enabled at runtime by setting an environment variable [^6] Set environment variable CUDACPP_RUNTIME_BLASCOLORSUM at runtime to use the cuBLAS (on Nvidia GPUs) or hipBLAS (on AMD GPUs) implementation of color sums.. Finally, as the BLAS library contains several switches targeting tensor cores, I also added another environment variable [^7] Set environment variable CUDACPP_RUNTIME_CUBLASTF32TENSOR at runtime to encourage cuBLAS to use tensor cores in color sums. This sets math mode CUBLAS_TF32_TENSOR_OP_MATH. to encourage BLAS to use tensor cores in the color sum (TF32 math mode).
@@
Figure 8: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, as a function of the number of events nevt processed in parallel in one GPU iteration. Lower is better. In each plot, four curves are shown, using the ihel3 (red) and ihel3p1 (black) codebases, in the CUDA kernel (solid) and CUDA/BLAS (dashed) implementations. The 3 plots correspond to 3 floating point precisions. The handling of the nGoodHel good helicities is different in the various scenarios. In the two CUDA kernel implementations, nGoodHel kernels are launched in separate helicity streams, each one on a GPU grid of size nevt: in particular, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementations: in ihel3b, nGoodHel separate BLAS color matrix multiplications are computed, in different helicity streams, using dual amplitude vectors with nevt events; in ihel3b, a single BLAS color matrix multiplications is computed, in the default stream, using dual amplitude vectors with nGoodHel*nevt events. Measurements performed for an NVidia V100 GPU, on the same node used to prepare Fig. 7.

### v3

@@ 3 Kernel splitting I: helicity streams, color sum
(1) Physics process $g\!\rightarrow\!t\bar{t} gg$$\rightarrow\!t\bar{t}g gg$$\rightarrow\!t\bar{t}gg gg$$\rightarrow\!t\bar{t}ggg gg$$\rightarrow\!t\bar{t}gggg gg$$g\!\rightarrow\!t\bar{t}ggggg$
(2) Feynman diagrams 3 16 123 1240 15495 231280
(3) SDE channels 3 15 105 945 N/A N/A
(4) Leading colors 2 6 24 120 720 5040
(5) Color matrix 2 x 2 6 x 6 24 x 24 120 x 120 720 x 720 5040 x 5040
(6) Helicities 16 32 64 128 256 512
(7) Wave functions 5 12 26 121 750 5834
(8) Code generation codebase ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p1
(9) Code generation mode mad mad mad mad sa sa
(10) Code generation d.p.g. any any any any $\leq$100 $\leq$1000
(11) Code generation d.p.f. any any any any $\geq$100 $\geq$1000
(12) Code generation time 2s 2s 4s 32s 13m 30h
(13) Feynman diagram groups 1 1 1 1 155 232
(14) Feynman diagram files 1 1 1 1 155 232
(15) LOC (color_sum.cc) 438 442 460 556 1156 5.4k
(16) LOC (diagramsXYZ.cc) 135 307 1986 27k 652k 18.7M
(17) Size (color_sum.cc) 24 kB 24 kB 28 kB 92 kB 2.6 MB 137 MB
(18) Size (diagramsXYZ.cc) 8 kB 16 kB 84 kB 1.0 MB 28.2 MB 721 MB
(19) Build time (five C++ modes) 6s 6s 6s 23s 3m 55m
(20) Build time (CUDA DCDIAG=0) 11s 11s 14s 6m 9m failed
(21) Build time (CUDA DCDIAG=1) 13s 13s 15s 3m 6m failed
@@
Table 2: Comparison of the computational complexity of the six physics processes considered in the tests described in this paper. The rows represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. of the diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for these six processes, this coincides with the number of good helicities with non-zero contributions to MEs); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the codebase used for code generation; (9) the code generation mode, “mad” for full generation including Fortran madevent and “sa” for standalone only; (9-10) the maximum number of diagrams per group (dpg) and the minimum number of diagrams per file (dpf) during code generation; (12) the code generation time on a typical CPU; (13-14) the resulting numbers of Feynman diagram groups and source code files; (15-17) the total number of lines of code and total size in the CUDAPP generated color_sum.cc and the sum of all diagramsXYZ.cc files; (19-21) the time for a parallel (-j) build of the software with a cold cache (CCACHE_RECACHE=1), either for all five C++ modes on an Intel Xeon Gold 6326, or for CUDA in each of two DCDIAG modes on an AMD EPYC 7313. All of these numbers are needed to describe the complexity of the ME calculation in a single standalone or madevent application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used, as of the ihel1 software. The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of ihel2, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel as of ihel2, and optionally via BLAS as of ihel3). The number of wavefunctions is particularly relevant to the ihel4-ihel6 software versions, which handle their storage and retrieval in GPU registers and/or GPU global memory in subtly different ways.
@@
As I briefly hinted above, one of the main aims of the new developments presented in this section is the fact that the ME calculation throughput (in MEs/s) on a GPU is generally quite low for small grids, i.e. small values of nevt, and only reaches a peak plateau for relatively large grids, i.e. large values of nevt. This is an issue that exists in CUDACPP since the very beginning of our developments: in Fig. 5 of our vCHEP2021 proceedings bib:vchep2021 ), for instance, we had shown that the ME throughput for the $gg\!\rightarrow\!t\bar{t}gg$ process only ramps up significantly with at least 16k events in the grid, and reaches the peak plateau even later, with 128k events in the grid. In the same paper, we had also already suggested that higher throughputs, possibly because of lower “register pressure”, might be achieved by splitting monolithic sigmaKin kernel into smaller kernels: two specific ideas that we had mentioned, in particular, were the possible use of different GPU threads to process different helicities in the same event, or the possible use of CUDA graphs to orchestrate a much larger number of smaller GPU kernels. I will come back to both of these ideas in the following.
@@
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this paper will mainly consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers that I had prepared for the HEP-SCORE benchmarking project bib:bmk ; it would be useful to repeat similar tests in the future, when updated HEP-SCORE containers are prepared using more recent versions of CUDACPP, notably those presented in this paper.
@@
In the following subsections, I will describe the first batch of three sets of kernel splitting changes that I developed, and which have already been integrated in a new production release of CUDACPP: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading, including some further refinements (ihel3p1). In the final subsection, I also describe a more recent set of color sum optimizations in the SIMD implementation (csm), which instead are not yet merged upstream. As shown in Fig. 1, these developments are sequential, e.g. ihel3 includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine for the ihel1, ihel2 and ihel3 scenarios is represented in Fig. 3, where they are also compared to the situation before any kernel splitting, ihel0. I will not provide new pseudo-code listings for these kernel splitting developments, but in some cases I will refer to the pseudo-code for version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, as well as for those in my second batch of changes described in the next section, are given in Fig. 4 for an NVidia V100 GPU at CERN, in Fig. 5 for an AMD MI200 GPU at LUMI, and in Fig. 6 for an Intel Xeon Silver 4216 CPU, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.
@@
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU at CERN, on a node equipped with Intel Xeon Silver 4216 CPUs. Code built using CUDA 12.0 and gcc 11.5. Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) cuBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, cuBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, cuBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, executed as a kernel at runtime (DCDIAG=0), without graphs.
@@
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU at LUMI, on a node equipped with AMD EPYC 7A53 CPUs. Code built using ROCm 6.0 and gcc 13.2. Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) hipBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, hipBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, hipBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime (DCDIAG=0), without graphs.
@@
Figure 6: Throughputs (ME/s) as a function of CUDACPP SIMD build mode for a single core of an Intel Xeon Silver 4216 CPU at CERN (using gcc 11.5). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel3p1) release v1.01.01; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, using code generated with a single diagram group.
@@ 3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
Figure 7: Time spent in the color sum as a fraction of the total spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 software. Lower is better. The 3 plots correspond to 3 floating point precisions. Measurements performed on an NVidia V100 GPU and a single core of an Intel Xeon Silver 4216 CPU, both on the same node at CERN (using CUDA 12.0 and gcc 11.5). The results for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
@@
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. The achieved performance is represented by the “ihel3” solid-red curve in Fig. 4. As mentioned above, these results are indistinguishable from those of ihel2, which I therefore decided not to show this explicitly to simplify the plot; the comparison, however, is available in a previous preprint of this article bib:preprintV1 .
@@
    i.e. as the sum of two quadratic forms $V^{t}(C)V$, where the real vector $V$ may represent either the real part $A$ or the imaginary part $B$ of the complex dual amplitude vector $J$. This decomposition is also used in GPU kernel color sums. Each calculation involves two steps. In mixed or float precision, for instance, the vector $(C)V$ is computed using cublasSgemm, while its dot product with $V^{t}$ is then computed using cublasSgemmStridedBatched. These two functions, as well as their double-precision and their HIP counterparts, are called through their abstractions via #define directives, in the header-only approach described in Ref. bib:chep2024 . The intermediate results of the first calculation, $(C)V$, are stored in GPU global memory using an additional buffer allocated in the MEK component (the allocation is done at runtime after determining the number of “good” helicities in each physics process, since the amount of memory allocated is proportional to the number of good helicities). As shown in the bottom-right diagram in Fig. 3, the BLAS calculations for different helicities are performed in separate GPU streams: technically, many BLAS handles are used, each associated to a different stream. Since the performance of the BLAS implementation of the color sum is generally worse than that using kernels, as discussed below, this is only available as an option, which must be explicitly enabled at runtime by setting an environment variable [^6] Set environment variable CUDACPP_RUNTIME_BLASCOLORSUM at runtime to use the cuBLAS (on Nvidia GPUs) or hipBLAS (on AMD GPUs) implementation of color sums.. Finally, as the BLAS library contains several switches targeting tensor cores, I also added another environment variable [^7] Set environment variable CUDACPP_RUNTIME_CUBLASTF32TENSOR at runtime to encourage cuBLAS to use tensor cores in color sums. This sets math mode CUBLAS_TF32_TENSOR_OP_MATH. to encourage BLAS to use tensor cores in the color sum (TF32 math mode).
@@
Figure 8: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, as a function of the number of events nevt processed in parallel in one GPU iteration. Lower is better. In each plot, four curves are shown, using the ihel3 (red) and ihel3p1 (black) codebases, in the CUDA kernel (solid) and CUDA/BLAS (dashed) implementations. The 3 plots correspond to 3 floating point precisions. The handling of the nGoodHel good helicities is different in the various scenarios. In the two CUDA kernel implementations, nGoodHel kernels are launched in separate helicity streams, each on a GPU grid of size nevt: in particular, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementations: in ihel3b, nGoodHel separate BLAS color matrix multiplications are computed, in different helicity streams, using dual amplitude vectors with nevt events; in ihel3p1b, a single BLAS color matrix multiplications is computed, in the default stream, using dual amplitude vectors with nGoodHel*nevt events. Measurements performed with an NVidia V100 GPU (using CUDA 12.0 and gcc 11.5), on the same node as in Fig. 7.